# Global-Decision-Focused Neural ODEs for Proactive Grid Resilience Management

### v2

@@ Abstract
Extreme hazard events such as wildfires and hurricanes increasingly threaten power systems, causing widespread outages and disrupting critical services. Recently, predict-then-optimize approaches have gained traction in grid operations, where system functionality forecasts are first generated and then used as inputs for downstream decision-making. However, this two-stage method often results in a misalignment between prediction and optimization objectives, leading to suboptimal resource allocation. To address this, we propose predict-all-then-optimize-globally (PATOG), a framework that integrates outage prediction with globally optimized interventions. At its core, our global-decision-focused (GDF) neural ODE model captures outage dynamics while optimizing resilience strategies in a decision-aware manner. Unlike conventional methods, our approach ensures spatially and temporally coherent decision-making, improving both predictive accuracy and operational efficiency. Experiments on synthetic and real-world datasets demonstrate significant improvements in outage prediction consistency and grid resilience.
Index Terms:
Outage prediction, proactive decision making, neural ODEs, decision-focused learning,
@@ I Introduction
Extreme hazard events such as wildfires, winter storms, hurricanes, and earthquakes can trigger widespread power outages, disrupting economic activity, threatening public safety, and complicating the delivery of critical services [1, 2]. For instance, the January 2025 wildfires in Los Angeles County, fueled by strong Santa Ana winds, led to the destruction of critical power infrastructure across vast areas, resulting in prolonged outages for over 400,000 customers [3]. Similarly, Hurricane Milton in October 2024 made landfall in Florida, bringing extreme winds and flooding that damaged the electrical grid and left more than 3 million homes and businesses without power for weeks [4], illustrating the catastrophic consequences of extreme weather on energy resilience.
@@
A fundamental challenge in strengthening power system resilience lies in the immense difficulty of restoring electricity after a severe event has caused widespread damage. The recent Los Angeles wildfires highlight how extreme winds can overwhelm firefighting efforts, making containment impossible even with abundant resources, let alone the rapidly repairing damaged transmission lines [5]. The aftermath of major storms presents similar challenges – damaged infrastructure can take weeks or even months to restore, leading to severe economic and social repercussions. However, the impact of such events can be significantly mitigated, and in some cases, entirely avoided through proactive resilience planning. For example, preemptive de-energization strategies have been successfully implemented in California to reduce wildfire risks [6], while grid hardening and strategic upgrades have improved system resilience in hurricane-prone regions [7].
@@
A widely used approach in such planning is the predict-then-optimize (PTO) framework, where forecasts inform downstream decisions [8]. However, conventional PTO approaches often fall short in high-stake decision making problems [9]. A high-quality forecast does not necessarily translate into effective decision-making, as small errors in predictive models can propagate into significant inefficiencies in downstream optimization tasks [10, 9]. This issue arises because conventional models are trained independently of decision-making objectives, meaning they may prioritize minimizing forecasting errors without considering how those errors impact critical decisions. As a result, even small deviations in predictions can lead to costly misallocations of resources, delayed responses, and suboptimal grid resilience strategies [11, 12]. Addressing this gap requires a more integrated approach that directly aligns predictive modeling with decision-making objectives, ensuring that forecasts are not only accurate but also actionable for proactive grid management.
@@
To address the challenges of proactive resilience planning, we first propose a new decision-making paradigm, referred to as predict-all-then-optimize-globally (PATOG), which integrates predictive outage modeling across all service units with a global decision-making process for grid resilience enhancement. To solve PATOG problems, we then introduce a global-decision-focused (GDF) neural ordinary differential equation (ODE) model, which simultaneously predicts the long-term evolution of system functionality (quantified by the number of power outages) and optimizes global resilience decisions. Inspired by epidemiological models, our method conceptualizes power outage progression within each service unit as a dynamical system, where the failure and restoration processes evolve based on transmission line conditions, local weather, and socioeconomic factors. These transition dynamics are parameterized using neural networks, allowing the model to adapt to different hazard scenarios and geographic conditions. Instead of treating each service unit independently, our method holistically forecasts the outage evolution across the power grid and derives a globally coordinated intervention strategy. This enables utilities to allocate resources more effectively, mitigating service disruptions in a spatially and temporally coherent manner.
To assess our framework, we conduct experiments on both real-world and synthetic datasets. The real dataset captures outage events from a 2018 Nor’easter in Massachusetts, combining county-level outage reports [13] with meteorological data from NOAA’s HRRR model [14] and socioeconomic indicators from the U.S. Census Bureau [15] (Fig. 1). The synthetic dataset simulates real outage propagation using a simplified SIR model, enabling controlled evaluation under varying weather conditions and grid configurations. These datasets provide a robust testbed for assessing GDF in realistic resilience planning scenarios. Our results demonstrate significant improvements in forecast consistency, decision efficiency, and overall grid resilience, showcasing the potential of our approach to enable smarter, more proactive energy infrastructure management in the face of increasing hazard-induced disruptions.
@@ I-A Related Work
Grid Operation Optimization. Grid optimization has been widely studied to improve power system resilience, covering areas such as distributed generator (DG) placement [16, 17, 18], infrastructure reinforcement [19], and dynamic power scheduling [20, 21]. However, traditional approaches often follow a two-stage predict-and-mitigate paradigm—first forecasting system conditions and then optimizing responses [22]. This disconnect between prediction and optimization results in suboptimal grid operations, particularly under high uncertainty, where even small forecasting errors can cause significant deviations from the optimal response [12, 11].
@@
Additionally, ordinary differential equations have been widely used to model dynamic systems, such as outage propagation, capturing evolving disruptions under various conditions [25, 26]. For instance, adaptations of the Susceptible-Infected-Recovered (SIR) model from epidemiology have been applied to simulate outage propagation, drawing parallels between power failures and disease spread [27].
@@
Decision-Focused Learning. Decision-focused learning (DFL) integrates predictive machine learning models with optimization, aligning training objectives with decision-making rather than purely maximizing predictive accuracy. Unlike traditional two-stage approaches, where predictions are first generated and then used as inputs for optimization, DFL enables end-to-end learning by backpropagating gradients through the optimization process. This is achieved via implicit differentiation of optimality conditions such as KKT constraints [28, 29] or fixed-point methods [30, 31]. For nondifferentiable optimizations, approximation techniques such as surrogate loss functions [8], finite differences [32], and noise perturbations [33] have been developed. Recent work has also explored integrating differential equation constraints directly into optimization models, enabling end-to-end gradient-based learning while ensuring compliance with system dynamic constraints [34, 35].
A well-studied class of DFL problems involves linear programs (LPs), where the Smart Predict-and-Optimize (SPO) framework [8] introduced a convex upper bound for gradient estimation, enabling cost-sensitive learning for optimization. Subsequent work has extended DFL to combinatorial settings, including mixed-integer programs (MIPs), using LP relaxations [36, 37]. Recent advances, such as decision-focused generative learning (Gen-DFL) [38], tackle the challenge of applying DFL in high-dimensional setting by using generative models to adaptively model uncertainty.
Differentiable Optimization. A key enabler of DFL is differentiable optimization (DO), which facilitates gradient propagation through differentiable optimization problems, aligning predictive models with decision-making objectives [39]. Recent advances extend DO to distributionally robust optimization (DRO) for handling uncertainty in worst-case scenarios, improving decision quality under data scarcity [40, 41]. Beyond predictive modeling, DO has advanced combinatorial and nonlinear optimization through implicit differentiation of KKT conditions [28], fixed-point methods [30], and gradient approximations via noise perturbation [33] and smoothing [32]. These techniques bridge forecasting with optimization, ensuring decision-aware learning. In this work, DO enables the backpropagation of resilience strategy losses, aligning the spatio-temporal outage prediction model with grid optimization objectives.
@@ II Predict All Then Optimize Globally
PATOG is particularly useful for grid resilience management, where decisions must consider complex interactions across all service units. A key example is the mobile generator distribution problem. In a conventional PTO setting, potential damage from an extreme weather event is first forecasted for each unit independently. Decisions, such as scheduling power generator deployments, are then made in isolation, without considering the evolving conditions of other units. This localized approach often leads to resource misallocation and suboptimal resilience outcomes. In contrast, PATOG embeds these interdependencies into a global optimization problem, enabling system-wide decision-making that improves predictive models by incorporating cross-unit interactions. This results in more robust and effective resilience planning.
@@ II-A Preliminaries: Decision-Focused Learning
The predict-then-optimize (PTO) has been extensively studied across a wide range of applications [22, 9]. It follows a two-step process: First, predicting the unknown parameters $\mathbf{c}$ using a model $f_{\theta}$ based on the input $\bm{z}$, denoted as $\hat{\bm{c}}\coloneqq f_{\theta}(\bm{z})$. Second, solving an optimization problem:
$\bm{x}^{*}(\hat{\mathbf{c}})=\arg\min_{\bm{x}}g(\bm{x},\hat{\mathbf{c}}),$ (1)
where $g$ is the objective function and $\bm{x}^{*}(\hat{\mathbf{c}})$ represents the optimal decision given the predicted parameters. This framework has numerous practical applications in power grid operations. For example, PTO is used in power grid operations to predict system stress using synchrophasor data, optimize outage management by forecasting disruptions and proactively dispatching restoration crews, and enhance renewable integration by predicting fluctuations in wind and solar generation to improve scheduling and grid balancing [22]. These predictive insights enable system operators to make informed strategic decisions, enhancing grid reliability and resilience.
However, the conventional two-stage approach does not always lead to high-quality decisions. In this approach, the model parameter $\theta$ is first trained to minimize a predictive loss, such as mean squared error. Then the predicted parameters $\hat{\mathbf{c}}$ are used to solve the downstream optimization. This separation between prediction and optimization can result in suboptimal decisions, as the prediction model is not directly optimized for decision quality [42].
To address this limitation, the decision-focused learning (DFL) integrates prediction with the downstream optimization process [9]. Instead of optimizing for predictive accuracy, DFL trains the model parameter $\theta$ by directly minimizing the decision regret [9]:
@@ II-B Proposed PATOG Framework
  • $U_{k}(t)\in\mathbb{Z}_{*}$: the number of unaffected customers;
  • $R_{k}(t)\in\mathbb{Z}_{*}$: the number of recovered customers.
@@
[…]
@@
where $g$ quantifies the decision loss of the action $\bm{x}$ based on the predicted future evolution of outage states $\mathbf{S}$. The transition function $f_{\theta}$ models the progression of unaffected, recovered, and outaged customers in each unit over time, influenced by both local weather conditions and socioeconomic factors, jointly represented as covariates $\bm{z}_{k}\in\mathbb{R}^{p}$. We note that most power outages during extreme weather stem from localized transmission line damage, with cascading failures being rare [43, 44]. Thus, we assume each unit evolves independently under its local conditions.
@@ III Global-Decision-Focused Neural ODEs
Figure 2: Overview of the proposed GDF framework. Given covariates $\bm{z}_{k}$ and initial states $\mathbf{S}_{k}(0)$ for all $K$ service units, a model parameterized by $\theta$ predicts the system states $\hat{\mathbf{S}}_{k}$ for all units. These predictions inform global decision-making, where optimal actions $\bm{x}^{*}(\hat{\mathbf{S}})$ minimize the global decision loss $g(\bm{x},\hat{\mathbf{S}})$. The framework optimizes $\theta$ by minimizing a global-decision-focused loss, regularized by a prediction-focused loss (e.g., MSE loss) to enhance predictive interpretability. Red arrows denote the backpropagation through $\nabla\ell_{\texttt{GDF}}(\theta)$, ensuring that the model learns both system-level decision quality and region-specific prediction accuracy.
@@ III-A Neural ODEs for Power Outages
Drawing inspiration from the Susceptible Infectious Recovered (SIR) models commonly used in epidemic modeling [45], we conceptualize power outages within a unit as the spread of a “virus”. In this analogy, outages propagate among customers due to local transmission line failures, while restorations provide lasting resistance to subsequent outages. We formalize this analogy with the following three assumptions: ($i$) The number of unaffected customers $U_{k}(t)$ decreases over time as some customers transition from being unaffected to experiencing outages. This transition is governed by the failure transmission rate, denoted by $\phi(\bm{z}_{k};\theta_{U})$. Inspired by epidemiological transmission rates, this rate quantifies how local conditions – such as weather patterns and other regional factors encapsulated by covariates $\bm{z}_{k}$ – influence the rate at which outages spread within the grid. ($ii$) Conversely, the number of customers with restored power ($R_{k}(t)$) gradually increases as the system operator repairs transmission lines and restores service. This process is captured by the restoration rate, denoted $\phi(\bm{z}_{k};\theta_{R})$. Both transmission and restoration rates, $\phi(\bm{z}_{k};\theta_{U})$ and $\phi(\bm{z}_{k};\theta_{R})$, are functions of the local covariates $\bm{z}_{k}$, and are modeled using deep neural networks. ($iii$) The total number of customers within each unit remains constant throughout the studied period. Based on these assumptions, we model the outage state transition for each unit $k$ in (3), i.e., $\mathrm{d}\mathbf{S}_{k}(t)/\mathrm{d}t$, as follows:
[…]
These ODEs capture the dynamic evolution of power outage states within each unit $k$. For notational simplicity, we use $\theta\coloneqq\{\theta_{U},\theta_{R}\}$ to denote their parameters jointly.
In practice, we work with discrete observations $\{t_{j}\}_{j=0}^{T}$ and adopt the Euler method to approximate the solution to the ODE model [26], i.e.,

### v3

@@ Abstract
Extreme hazard events such as wildfires and hurricanes increasingly threaten power systems, causing widespread outages and disrupting critical services. To mitigate these risks, grid operators must act proactively—pre-positioning crews and resources or scheduling hardening activities—guided by forecasts that directly support operational objectives. Traditionally, utilities and agencies have followed a predict-then-optimize paradigm: first generating impact forecasts, then using them to inform response planning. However, this two-step approach inherently separates forecasting from decision-making, overlooking how forecast errors propagate into downstream actions and often producing misaligned or suboptimal plans. We address this gap with predict-all-then-optimize-globally (PATOG), a framework that unifies outage prediction with globally optimized interventions. At its core, our global-decision-focused (GDF) Neural ODE model jointly captures outage dynamics and optimizes pre-event resilience strategies in a decision-aware manner. Unlike conventional methods, our approach ensures spatially and temporally coherent decisions, enhancing both predictive accuracy and operational efficiency. Evaluations on synthetic and real-world data show that GDF reduces decision regret by up to 75% compared to two-stage baselines, enabling more effective proactive planning.
@@ I Introduction
Extreme hazard events such as wildfires, winter storms, hurricanes, and earthquakes can trigger widespread power outages, disrupting economic activity, threatening public safety, and complicating the delivery of critical services [Handmer2012, Kenward2014]. For instance, the January 2025 wildfires in Los Angeles County, fueled by strong Santa Ana winds, led to the destruction of critical power infrastructure across vast areas, resulting in prolonged outages for over 400,000 customers [reuters2024hurricane]. Similarly, Hurricane Milton in October 2024 made landfall in Florida, bringing extreme winds and flooding that damaged the electrical grid and left more than 3 million homes and businesses without power for weeks [reuters2025wildfires], illustrating the catastrophic consequences of extreme weather on energy resilience.
@@
A fundamental challenge in strengthening power system resilience lies in the immense difficulty of restoring electricity after a severe event has caused widespread damage. The recent Los Angeles wildfires highlight how extreme winds can overwhelm firefighting efforts, making containment impossible even with abundant resources, let alone the rapidly repairing damaged transmission lines [CNN2025Wildfires]. The aftermath of major storms presents similar challenges—damaged infrastructure can take weeks or even months to restore, leading to severe economic and social repercussions. However, the impact of such events can be significantly mitigated, and in some cases, entirely avoided through proactive resilience planning. For example, preemptive de-energization strategies have been successfully implemented in California to reduce wildfire risks [wiki:2019_CA_power_shutoffs], while grid hardening and strategic upgrades have improved system resilience in hurricane-prone regions [HoustonChronicle2025Grid].
@@
A widely used approach in such planning is the predict-then-optimize (PTO) framework, where forecasts inform downstream decisions [elmachtoub_smart_2022]. However, conventional PTO approaches often fall short in high-stake decision making problems [mandi_decision-focused_2024]. A high-quality forecast does not necessarily translate into effective decision-making, as small errors in predictive models can propagate into significant inefficiencies in downstream optimization tasks [fernandez2022causal, mandi_decision-focused_2024]. This issue arises because conventional models are trained independently of decision-making objectives, meaning they may prioritize minimizing forecasting errors without considering how those errors impact critical decisions. As a result, even small deviations in predictions can lead to costly misallocations of resources, delayed responses, and suboptimal grid resilience strategies [10855832, 10711393]. Addressing this gap requires a more integrated approach that directly aligns predictive modeling with decision-making objectives, ensuring that forecasts are not only accurate but also actionable for proactive grid management.
@@
To address the challenges of proactive resilience planning, we first propose a new decision-making paradigm, referred to as predict-all-then-optimize-globally (PATOG), which integrates predictive outage modeling across all service units with a global decision-making process for grid resilience enhancement. Examples of such planning problems for better pre-event preparation include scheduling dispatch of mobile generators and planning budget-constrained power line undergrounding. To solve PATOG problems, we introduce a global-decision-focused (GDF) neural ordinary differential equation (ODE) model, which simultaneously predicts the long-term evolution of system functionality (quantified by the number of power outages) and optimizes global resilience decisions. Inspired by epidemiological models, our method conceptualizes power outage progression within each service unit as a dynamical system, where the failure and restoration processes evolve based on transmission line conditions, local weather, and socioeconomic factors. These transition dynamics are parameterized using neural networks, allowing the model to adapt to different hazard scenarios and geographic conditions. Instead of treating each service unit independently, our method holistically forecasts the outage evolution across the power grid and derives a globally coordinated intervention strategy. This enables utilities to allocate resources more effectively, mitigating service disruptions in a spatially and temporally coherent manner.
To assess our framework, we conduct experiments on both real-world and synthetic datasets. The real dataset captures outage events from a 2018 Nor’easter in Massachusetts, combining county-level outage reports [maema2020power] with meteorological data from NOAA’s HRRR model [noaa_hrrr] and socioeconomic indicators from the U.S. Census Bureau [USCensusBureau2017ACS] (Fig. 1). The synthetic dataset simulates real outage propagation using a simplified SIR model, enabling controlled evaluation under varying weather conditions and grid configurations. These datasets provide a robust testbed for assessing GDF in realistic resilience planning scenarios. Our results demonstrate significant improvements in forecast consistency, decision efficiency, and overall grid resilience, showcasing the potential of our approach to enable smarter, more proactive energy infrastructure management in the face of increasing hazard-induced disruptions.
@@ I-A Related Work
Grid Operation Optimization. Grid optimization has been widely studied to improve power system resilience, covering areas such as distributed generator (DG) placement [8668633, QIN2024752, 9117454], infrastructure reinforcement [8273985], and dynamic power scheduling [5963580, 990600]. However, traditional approaches often follow a two-stage predict-and-mitigate paradigm—first forecasting system conditions and then optimizing responses [7080837]. This disconnect between prediction and optimization results in suboptimal grid operations, particularly under high uncertainty, where even small forecasting errors can cause significant deviations from the optimal response [10711393, 10855832].
@@
Additionally, ordinary differential equations have been widely used to model dynamic systems, such as outage propagation, capturing evolving disruptions under various conditions [zhang2024recurrent, chen2018neural]. For instance, adaptations of the Susceptible-Infected-Recovered (SIR) model from epidemiology have been applied to simulate outage propagation, drawing parallels between power failures and disease spread [9646114].
@@
Decision-Focused Learning. Decision-focused learning (DFL) integrates predictive machine learning models with optimization, aligning training objectives with decision-making rather than purely maximizing predictive accuracy. Unlike traditional two-stage approaches, where predictions are first generated and then used as inputs for optimization, DFL enables end-to-end learning by backpropagating gradients through the optimization process. This is achieved via implicit differentiation of optimality conditions such as KKT constraints [amos2017optnet, gould2021deep] or fixed-point methods [kotary2023folded, wilder2019end]. For nondifferentiable optimizations, approximation techniques such as surrogate loss functions [elmachtoub_smart_2022], finite differences [vlastelica2019differentiation], and noise perturbations [berthet2020learning] have been developed. Recent work has also explored integrating differential equation constraints directly into optimization models, enabling end-to-end gradient-based learning while ensuring compliance with system dynamic constraints [Fioretto:ICLR25, jacquillat2024branch].
A well-studied class of DFL problems involves linear programs (LPs), where the Smart Predict-and-Optimize (SPO) framework [elmachtoub_smart_2022] introduced a convex upper bound for gradient estimation, enabling cost-sensitive learning for optimization. Subsequent work has extended DFL to combinatorial settings, including mixed-integer programs (MIPs), using LP relaxations [mandi2020smart, wilder2019melding]. Recent advances, such as decision-focused generative learning (Gen-DFL) [wang2025gendfldecisionfocusedgenerativelearning], tackle the challenge of applying DFL in high-dimensional setting by using generative models to adaptively model uncertainty.
Differentiable Optimization. A key enabler of DFL is differentiable optimization (DO), which facilitates gradient propagation through differentiable optimization problems, aligning predictive models with decision-making objectives [cvx]. Recent advances extend DO to distributionally robust optimization (DRO) for handling uncertainty in worst-case scenarios, improving decision quality under data scarcity [zhu2022distributionally, chen2025uncertainty]. Beyond predictive modeling, DO has advanced combinatorial and nonlinear optimization through implicit differentiation of KKT conditions [amos2017optnet], fixed-point methods [kotary2023folded], and gradient approximations via noise perturbation [berthet2020learning] and smoothing [vlastelica2019differentiation]. These techniques bridge forecasting with optimization, ensuring decision-aware learning. In this work, DO enables the backpropagation of resilience strategy losses, aligning the spatio-temporal outage prediction model with grid optimization objectives.
@@ II Predict All Then Optimize Globally
PATOG is particularly useful for grid resilience management, where decisions must consider complex interactions across all service units. A key example is the mobile generator deployment problem. In a conventional PTO setting, potential damage from an extreme weather event is first forecasted for each unit independently. Decisions, such as scheduling power generator deployments, are then made in isolation, without considering the evolving conditions of other units. This localized approach often leads to resource misallocation and suboptimal resilience outcomes. In contrast, PATOG embeds these interdependencies into a global optimization problem, enabling system-wide decision-making that improves predictive models by incorporating cross-unit interactions. This results in more robust and effective resilience planning.
@@ II-A Preliminaries: Decision-Focused Learning
The predict-then-optimize (PTO) has been extensively studied across a wide range of applications [7080837, mandi_decision-focused_2024]. It follows a two-step process: First, predicting the unknown parameters $\mathbf{c}$ using a model $f_{\theta}$ based on the input $\bm{z}$, denoted as $\hat{\bm{c}}\coloneqq f_{\theta}(\bm{z})$. Second, solving an optimization problem:
[…]
However, the conventional two-stage approach does not always lead to high-quality decisions. In this approach, the model parameter $\theta$ is first trained to minimize a predictive loss, such as mean squared error. Then the predicted parameters $\hat{\mathbf{c}}$ are used to solve the downstream optimization. This separation between prediction and optimization can result in suboptimal decisions, as the prediction model is not directly optimized for decision quality [kotary2021end].
To address this limitation, the decision-focused learning (DFL) integrates prediction with the downstream optimization process [mandi_decision-focused_2024]. Instead of optimizing for predictive accuracy, DFL trains the model parameter $\theta$ by directly minimizing the decision regret [mandi_decision-focused_2024]:
@@ II-B Proposed PATOG Framework
  • $R_{k}(t)\in\mathbb{Z}_{*}$: the number of restored customers.
@@
$\text{s.t.}\quad\frac{\mathrm{d}\mathbf{S}_{k}(t)}{\mathrm{d}t}=$ $~f_{\theta}(\mathbf{S}_{k}(t),\bm{z}_{k}),~\forall k,$ (3)
@@
where $g$ quantifies the decision loss of the action $\bm{x}$ based on the predicted future evolution of outage states $\mathbf{S}$. The transition function $f_{\theta}$ models the progression of unaffected, restored, and outaged customers in each unit over time, influenced by both local weather conditions and socioeconomic factors, jointly represented as covariates $\bm{z}_{k}\in\mathbb{R}^{p}$. We note that most power outages during extreme weather stem from localized transmission line damage, with cascading failures being rare [zhu2021quantifying, 10479971]. Thus, we assume each unit evolves independently under its local conditions.
@@ III Global-Decision-Focused Neural ODEs
[…]
@@ III-A Neural ODEs for Power Outages
Drawing inspiration from the Susceptible Infectious Recovered (SIR) models commonly used in epidemic modeling [kosma2023neural], we conceptualize power outages within a unit as the spread of a “virus”. In this analogy, outages propagate among customers due to local transmission line failures, while restorations provide lasting resistance to subsequent outages. We formalize this analogy with the following three assumptions: ($i$) The number of unaffected customers $U_{k}(t)$ decreases over time as some customers transition from being unaffected to experiencing outages. This transition is governed by the failure transmission rate, denoted by $\phi(\bm{z}_{k};\theta_{U})$. Inspired by epidemiological transmission rates, this rate quantifies how local conditions—such as weather patterns and other regional factors encapsulated by covariates $\bm{z}_{k}$—influence the rate at which outages spread within the grid. ($ii$) Conversely, the number of customers with restored power ($R_{k}(t)$) gradually increases as the system operator repairs transmission lines and restores service. This process is captured by the restoration rate, denoted $\phi(\bm{z}_{k};\theta_{R})$. Both transmission and restoration rates, $\phi(\bm{z}_{k};\theta_{U})$ and $\phi(\bm{z}_{k};\theta_{R})$, are functions of the local covariates $\bm{z}_{k}$, and are modeled using deep neural networks. (iiiiii) The total number of customers within each unit remains constant throughout the studied period. Based on these assumptions, we model the outage state transition for each unit $k$ in (3), i.e., $\mathrm{d}\mathbf{S}_{k}(t)/\mathrm{d}t$, as follows:
$\begin{cases}\begin{aligned} \mathrm{d}U_{k}(t)/\mathrm{d}t&=-\phi(\bm{z}_{k};\theta_{U})Y_{k}(t)U_{k}(t),\\ \mathrm{d}R_{k}(t)/\mathrm{d}t&=\phi(\bm{z}_{k};\theta_{R})Y_{k}(t),\\ \mathrm{d}Y_{k}(t)/\mathrm{d}t&=-\mathrm{d}U_{k}(t)/\mathrm{d}t-\mathrm{d}R_{k}(t)/\mathrm{d}t.\end{aligned}\end{cases}$ (5)
These ODEs capture the dynamic evolution of power outage states within each unit $k$. For notational simplicity, we use $\theta\coloneqq\{\theta_{U},\theta_{R}\}$ to denote their parameters jointly. This representation is particularly suitable for modeling weather-driven outage events with evolving spatio-temporal dynamics, such as those caused by winter storms or hurricanes, where the covariate-dependent rates $\phi(\bm{z}_{k};\theta_{U})$ and $\phi(\bm{z}_{k};\theta_{R})$ can capture exogenous meteorological forces leading to the hazard events.
In practice, we work with discrete observations $\{t_{j}\}_{j=0}^{T}$ and adopt the Euler method to approximate the solution to the ODE model [chen2018neural], i.e.,